girl is 5 times older than her brother, in 6 years she will be twice as old as her brother. How old are they now?
Answer by stanbon(75887)

girl is 5 times older than her brother, in 6 years she will be twice as old as her brother. How old are they now?
----
Equations:
g = 5b
(g+6) = 2(b+6)
----
Substitute for "g" and solve for "b":
(5b+6) = 2b+12
3b = 6
b = 2 (brother's age now)
---
Solve for "g":
g = 5b
g = 10 (girl's age now)
==========================
Cheers,
Stan H.
=============